Report #38 โ Submitted 2026-06-22 06:55:55
| No | Observation | Status |
|---|---|---|
| โ No observations recorded โ | ||
| Furnace | Mass Received (Ton) | Consumed (Ton) | Metal Added (Ton) | Mass End of Shift (Ton) | Remarks |
|---|---|---|---|---|---|
| HA
min: 28T
|
28.00 | 6.12 | 0.00 | 21.88 | we found havy dross inside the HA so we done the cleaning prosses |
| HC
min: 24T
|
14.80 โ Below minimum
|
0.00 | 0.00 | 14.80 | no pump |
| SWM1 | 0.00 | 0.00 | 0.00 | 0.00 | โ |
| SWM2 | 0.00 | 0.00 | 0.00 | 0.00 | โ |
| Stage | Condition | Comments |
|---|---|---|
| Received | Bad | work is going on for repair from HA to HC half complete From HA to SWM2 not start yet |
| Hand Over | Bad | work is going on for repair from HA to HC half complete From HA to SWM2 not start yet -- extra launder fix to transfer the metal for taping |
| M/C No. | Status | Mold Type | NDT OK | HC & FC Reject | NDT Reject | Total | Rejection % | H&F Rej % | NDT Rej % | Remarks |
|---|---|---|---|---|---|---|---|---|---|---|
| 5 | Running | 640-15" | 26 | 5 | 2 | 33 | 21.2% | 15.2% | 6.1% | stop at 10 PM |
| 6 | Running | 640-17"-m2 | 52 | 7 | 2 | 61 | 14.8% | 11.5% | 3.3% | stop at 10:00pm -- start again 4:30am |
| 7 | Stopped | 640-17-m1 | 60 | 7 | 1 | 68 | 11.8% | 10.3% | 1.5% | stop at 10:00pm -- start again 2:30 am |
| 14 | Stopped | r62-8" | 31 | 4 | 1 | 36 | 13.9% | 11.1% | 2.8% | stop at 10 PM |
| 16 | Running | 640-16"-m2 | 61 | 7 | 1 | 69 | 11.6% | 10.1% | 1.5% | stop at 10:00pm -- start again 3:30am |
| TOTAL PRODUCTION | 267 | โ | โ | โ | โ | |||||
| TOTAL REJECTION | 37 | 13.9% | 11.2% | 2.6% | โ | |||||
| M/C | Issue | DownTime (min) | Action Taken | Pending Action | Responsible | Status | Odoo Notification |
|---|---|---|---|---|---|---|---|
| laser mark 2 | scaner | 0 min | maintenance team | no | maintenance team | Closed | yes |
| prehating machine no 1 | flaim not working contusions | 0 min | maintenance team | yes | maintenance team | In Progress | yes |
| LPDC 4 | air pressure hose disconnected | 0 min | maintenance team inform | no | maintenance team | Closed | yes |
| Die Model | Status | Remarks |
|---|---|---|
| R69-18"-M1 | Completed | full service |
| 640-15"-M1 | Completed | upper mold + 4 side molds complete |
While Shaft 5 was running, we noticed that the Holding Furnace (HA) temperature was not reaching the required 800ยฐC. The actual temperature was around 726ยฐC, and after each ladle transfer, the temperature dropped further to approximately 640ยฐC or below. We inspected the HA and found a heavy buildup of dross inside the furnace. The pump was lifted to allow cleaning operations. During the inspection, we found that the impeller was not rotating. We attempted to rotate it manually, but it was completely stuck. At that stage, we suspected either a mechanical failure of the impeller or blockage caused by dross accumulation. After lifting the assembly further, we discovered a piece of metal scrap lodged in the impeller, preventing it from rotating. The scrap may have entered the area during adding element operations or while removing dross from another location inside the furnace. The cleaning was completed, the obstruction was removed, and the pump was reinstalled. Melting operations were restarted at approximately 12:30 PM. As a result of this issue, all casting machines were stopped due to insufficient metal temperature. Even after cleaning, the temperature of the first ladle remained below the required casting temperature. Therefore, the first ladle was blocked and the second ladle was used for casting operations. Root Cause (Suspected): Metal scrap became trapped in the impeller, causing it to stop rotating. Reduced circulation inside the Holding Furnace led to low metal temperature and significant temperature drops during ladle transfers. Impact: Holding Furnace temperature could not reach the required 800ยฐC. Casting operations were interrupted. All casting machines were stopped until corrective actions were completed. Corrective Action: Pump lifted and furnace cleaned. Metal scrap removed from the impeller. Pump reinstalled and melting operations resumed. First low-temperature ladle blocked and the second ladle used for production.